# Entanglement, defects, and $T\bar{T}$ on a black hole background

### v1

@@ 1 Introduction
The black hole information paradox has been an active research topic in semi-classical and quantum gravity for several decades. In recent developments, a possible resolution to this paradox was proposed which considered a toy model of a quantum field theory coupled to a semi-classical gravity, finally reproducing the infamous Page curve [1, 2, 3] for the Hawking radiation. The aptly named island formalism involves the emergence of spacetime regions called ”islands” at late times in the entanglement wedge of the radiation bath, leading to the purification of the outgoing Hawking quanta [4, 5, 6, 7, 8, 9]. Inspired by the quantum extremal surface (QES) prescription [10], the island formula was proposed for computing the fine-grained entanglement entropy of the Hawking radiation.
@@
The double holographic prescription [4, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] provides an interesting alternative perspective on the island formalism. It described a $d+1$-dimensional gravitational theory in a braneworld as a dual to a $d$-dimensional conformal field theory (CFTd) coupled to semi-classical gravity on the End-of-the-world (EOW) brane. This coupled description, referred to as the brane or the lower-dimensional effective description, may be obtained via a partial reduction of the bulk Anti-de Sitter (AdS) geometry [22, 23, 24, 25, 26, 27]. Furthermore, this gravity-matter system on the EOW brane is the holographic dual to a quantum mechanical system located at the boundary of the CFTd, known as the boundary description. The calculation of entanglement entropy in this context aligns with the Ryu-Takayanagi (RT) prescription in the bulk AdSd+1 geometry.
@@
On a separate note, the author in [28] introduced the boundary conformal field theory (BCFT) as a CFT defined on a domain with a boundary. The AdS/BCFT correspondence was subsequently described in [29] and was further developed in recent years (see [30, 11, 31, 32, 33, 34, 35, 36]). The AdS3/BCFT2 correspondence describes an asymptotically AdS3 geometry truncated by a codimension-1 EOW brane with Neumann boundary conditions as the holographic dual for a BCFT2. Due to modified homology conditions of the RT surfaces in this framework, the holographic entanglement entropy formula involves extremal surfaces that terminate on the EOW brane [29, 30].
@@
This AdS3/BCFT2 scenario was further developed in [22], where the authors considered the bulk as a defect spacetime with conformal matter on the EOW brane. In this case, the gravity region relevant to determining the QES is localized only on the EOW brane and emerges via a partial Randall-Sundrum reduction of the bulk geometry. In the lower-dimensional effective description, transparent boundary conditions were imposed at the interface of the gravitational and non-gravitational regions of the CFT2. This description, known as the defect extremal surface (DES) prescription, was proposed as a holographic counterpart of the island formula. Computations of the fine-grained entanglement entropy using the DES formula and the boundary QES prescription were shown to agree. Subsequently, the authors in [23] employed the DES formula to compute the entanglement entropy for subsystems in radiation baths associated with a $2d$ eternal black hole.
@@
To assess the robustness of the island framework and its double holographic generalizations, it is crucial to examine them in a less conventional gravitational setting. The solvable nature of CFTs deformed by an irrelevant operator formed from the stress-energy tensors [37, 38, 39] creates a suitable environment for this purpose. These irrelevant deformations, which are extensively studied in the literature [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 27, 66, 67, 68, 69], introduce a finite cut-off in the AdS3 geometries [70] with Dirichlet boundary conditions. [^1] Note that alternate holographic proposals for $T\bar{T}$ deformed theories, such as [71, 41, 72, 73, 74, 75], exist in the literature and are expected to coincide under certain limits.
@@
In [76, 77], the authors computed the entanglement entropy for bipartite pure states using both the island and the DES formula within a cut-off AdS3 spacetime truncated by an EOW brane. This setup involves two boundaries- an EOW brane characterizing the AdS/BCFT correspondence, and a finite cut-off surface that arises due to $T\bar{T}$ deformed theories. These surfaces differ fundamentally due to the distinct boundary conditions imposed and introduce subtleties in the standard AdS/BCFT correspondence, which require careful analysis. In the boundary description, this construction is equivalent to a $T\bar{T}$ deformed BCFT2 obtained via partial dimensional reduction of the braneworld gravity glued to a non-gravitating CFT with $T\bar{T}$ deformation [76].
@@
In [22], the proposed duality between the island formula and the DES prescription has been verified for simple braneworld models where the induced geometry on the EOW brane was Poincaré AdS2 and the BCFT was on a flat and non-gravitating background. However, it is crucial to verify this duality in more general settings where the induced metric on the brane and the BCFT background might not be so simple. In [78] the authors considered a gravitational FRW cosmology on the EOW brane coupled to a flat CFT on the asymptotic boundary, and verified the equivalence between the island and the DES prescriptions using the entanglement entropy.
@@
In this article, we investigate the proposed duality between the island and the DES prescription for braneworld models where the radiation bath is not on a flat background. In particular, we consider the model described in [20, 79], where the authors consider a bulk AdS3 black string geometry truncated by a Karch-Randall (KR) brane, dual to a BCFT2 considered on an eternal AdS2 black hole background. In this model, the metric induced on the brane is also an eternal AdS2 black hole. We modify this model to include conformal matter localized on the EOW brane, and utilizing the island and the DES prescription, we compute the entanglement entropy for subsystems considered in radiation baths associated with an eternal AdS2 black hole. We show that the results obtained using the DES prescription align with those obtained from the island formula. We further extend our analysis to an AdS3 black string geometry with a finite cut-off arising from the $T\bar{T}$ deformation and truncated by the EOW brane, and compute the entanglement entropy up to first order corrections for subsystems in the radiation baths. Once again, we demonstrate an agreement between the results obtained using the bulk DES and the island prescriptions.
@@
The remainder of the article is as follows. In section 2 we briefly discuss some necessary concepts, such as the Defect AdS3/BCFT2, the island and the DES formula for the entanglement entropy, the braneworld model of a holographic BCFT2 located on an AdS2 black hole background, and $T\bar{T}$ deformation. In section 3, we discuss the computations of the entanglement entropy for subsystems in the radiation baths in this braneworld model, utilizing both the island and the DES prescription. Subsequently, we determine the correction to the entanglement entropy in the cut-off AdS3 black string model arising from $T\bar{T}$ deformation truncated by the EOW brane in section 4. We generate and compare Page curves for both the undeformed and the $T\bar{T}$ deformed scenarios in section 5, and finally, we conclude by summarizing our results and discussing their implications in section 6.
@@ 2.1 Defect AdS3/BCFT2
In this section we briefly review the AdS3/BCFT2 correspondence first proposed in [29, 30]. A boundary conformal field theory (BCFT) is defined as a conformal field theory (CFT) with a boundary, where conformal boundary conditions are applied. The holographic dual of a BCFT [28] is an asymptotically AdS spacetime truncated by a co-dimension 1 end-of-the-world (EOW) brane where Neumann boundary conditions are applicable. The bulk action in gives as [29, 30]
$I_{0}=-\frac{1}{16\pi G_{N}}\int_{N}\sqrt{g}(R-2\Lambda)-\frac{1}{8\pi G_{N}}\int_{Q}\sqrt{h}(K-(d-1)T),$ (2.1)
@@
$K_{ab}=(K-T)h_{ab}.$ (2.2)
Figure 1: Holographic dual of a BCFT2 defined on a half plane $(x>0)$. The EOW brane is considered to be at a constant hyperbolic angle $\rho_{0}$. (Figure modified from [22].)
@@
$ds^{2}=d\rho^{2}+\cosh^{2}\rho\left(\frac{d\tau^{2}+dy^{2}}{y^{2}}\right),$ (2.3)
@@
[…]
@@
$K_{ab}=\tanh\rho_{0}\,\ h_{ab},\qquad T=\tanh\rho_{0}.$ (2.5)
This AdS3/BCFT2 setup was further developed in [22, 23] where the authors incorporated conformal matter localised on a tensionless EOW brane. This induces a tension on the brane, which result in the brane to relocate to some constant angle dependent on the tension. The bulk action in this modified setup is given by
$I=I_{0}+\int_{Q}\sqrt{h}\mathcal{L}_{Q},$ (2.6)
where $\mathcal{L}_{Q}$ is the Lagrangian corresponding to the conformal matter on the brane. [^2] See [13] for some examples. Once again by variation of this modified action with respect to the induced metric $h_{ab}$ and implementation of the Neumann boundary condition we arrive at [29, 22, 23]
$K_{ab}=(K-T)h_{ab}+8\pi G_{N}\langle T_{ab}\rangle,\qquad\langle T_{ab}\rangle=\frac{2}{\sqrt{h}}\frac{\partial\mathcal{L}_{Q}}{\partial h_{ab}}.$ (2.7)
@@ 2.2 Entanglement Entropy and the DES formula
In the modified scenario of defect AdS3/BCFT2 discussed in the previous subsection, the entanglement entropy of a subsystem $A$ undergoes modifications to incorporate contributions from the conformal matter localised on the brane. In this framework the fine grained entanglement entropy of a subsystem $A$ defined in the radiation bath is computed using the island formula described in [4, 80, 5] as
$S_{Is}=\text{min}\,\ \text{ext}_{X}\Big\{S_{eff}(A\cup I_{S}(A))+S_{area}(X)\Big\},\qquad X=\partial I_{S}(A)$ (2.8)
@@
From the holographic perspective, the Defect Extremal Surface (DES) formula was introduced as a holographic counterpart of the island formula in [22], where the authors enhanced the Quantum Extremal Surface (QES) formula to incorporate contributions from the bulk defect matter localized on the EOW brane. This leads to the modification of the Ryu-Takayanagi (RT) formula for the holographic entanglement entropy to the DES formula as
$S_{DES}=\text{min}\,\ \text{ext}_{\Gamma_{A},X}\Big\{S_{RT}(\Gamma_{A})+S_{defect}(D)\Big\},\qquad X=\Gamma_{A}\cap D,$ (2.9)
@@ 2.3 Holographic BCFT2 in a Black Hole background
In [20, 79], the authors examined the entanglement entropy of subsystems considered on the asymptotic boundary of an AdS3 black string geometry truncated by a Karch-Randall (KR) brane. The bulk geometry is described by the metric
$ds^{2}=\frac{\cosh^{2}\rho}{u^{2}}\bigg[-\left(1-\frac{u}{u_{h}}\right)dt^{2}+\frac{du^{2}}{\left(1-\frac{u}{u_{h}}\right)}\bigg]+d\rho^{2},$ (2.10)
@@
$X_{2}=\frac{2\sqrt{u_{h}^{2}-uu_{h}}}{u}\cosh\frac{2\pi t}{\beta}\cosh\rho,\qquad X_{3}=\sinh\rho,$ (2.11)
@@
$ds^{2}=\eta_{AB}dX^{A}dX^{B},\qquad\eta_{AB}=\text{diag}(-1,-1,1,1),$ (2.12)
@@
$X_{A}X^{A}=-1.$ (2.13)
Here $\beta=4\pi u_{h}$ is the inverse Hawking temperature. Using the above embedding relations, the holographic entanglement entropy for radiation subsystems considered on both the left and right copies of the TFD state corresponding to an eternal black hole was evaluated in [79] as
$S_{A}=\begin{cases}\frac{c}{3}(\rho_{\epsilon}+\rho_{b})&\text{Boundary channel,}\\ \frac{c}{6}\left(\log\left[\frac{u_{h}}{LR}\left(\Delta_{L}+\Delta_{R}+2\sqrt{\Delta_{L}\Delta_{R}}\cosh\frac{t}{u_{h}}\right)\right]+2\rho_{\epsilon}\right)\qquad&\text{Bulk channel,}\end{cases}$ (2.14)
where $\Delta_{L}=u_{h}-L$, $\Delta_{R}=u_{h}-R$, and $\rho_{\epsilon}$ is related to the UV cut-off $\epsilon$ as $\rho_{\epsilon}=\log\frac{2}{\epsilon}$. These results were also reproduced from the boundary description using BCFT correlators of twist fields in a black hole background in [79].

### v2

@@ Abstract
Keywords:
AdS/CFT Correspondence, Black Holes, Defects
arxiv: 2511.05256
@@ 1 Introduction
The black hole information paradox has been an active research topic in semi-classical and quantum gravity for several decades. In recent developments, a possible resolution to this paradox was proposed which considered a toy model of a quantum field theory coupled to a semi-classical gravity, finally reproducing the infamous Page curve Page:1993wv ; Page:1993df ; Page:2013dx for the Hawking radiation. The aptly named island formalism involves the emergence of spacetime regions called ’islands’ at late times in the entanglement wedge of the radiation bath, leading to the purification of the outgoing Hawking quanta Almheiri:2019hni ; Almheiri:2019psf ; Almheiri:2019qdq ; Almheiri:2019psy ; Almheiri:2019yqk ; Almheiri:2020cfm ; Geng:2025rov ; Penington:2019npb . Inspired by the quantum extremal surface (QES) prescription Engelhardt:2014gca , the island formula was proposed for computing the fine-grained entanglement entropy of the Hawking radiation.
@@
The double holographic prescription Almheiri:2019hni ; Sully:2020pza ; Rozali:2019day ; Chen:2020uac ; Chen:2020hmv ; Grimaldi:2022suv ; Suzuki:2022xwv ; Geng:2020qvw ; Geng:2020fxl ; Geng:2021iyq ; Geng:2021mic ; Geng:2021hlu ; Jain:2023xta ; Geng:2024xpj provides an interesting alternative perspective on the island formalism. It described a $d+1$-dimensional gravitational theory in a braneworld as a dual to a $d$-dimensional conformal field theory (CFTd) coupled to semi-classical gravity on the End-of-the-world (EOW) brane. This coupled description, referred to as the brane or the lower-dimensional effective description, may be obtained via a partial reduction of the bulk Anti-de Sitter (AdS) geometry Deng:2020ent ; Chu:2021gdb ; Li:2021dmf ; Basu:2022reu ; Shao:2022wrm ; Basu:2024xjq . Furthermore, this gravity-matter system on the EOW brane is the holographic dual to a quantum mechanical system located at the boundary of the CFTd, known as the boundary description. The calculation of entanglement entropy in this context aligns with the Ryu-Takayanagi (RT) prescription in the bulk AdSd+1 geometry.
@@
On a separate note, the author in Cardy:2004hm introduced the boundary conformal field theory (BCFT) as a CFT defined on a domain with a boundary. The AdS/BCFT correspondence was subsequently described in Takayanagi:2011zk and was further developed in recent years (see Fujita:2011fp ; Sully:2020pza ; Kastikainen:2021ybu ; Izumi:2022opi ; Cavalcanti:2018pta ; Magan:2014dwa ; Cavalcanti:2020rsp ; Takayanagi:2020njm ; Geng:2025efs ). The AdS3/BCFT2 correspondence describes an asymptotically AdS3 geometry truncated by a codimension-1 EOW brane with Neumann boundary conditions as the holographic dual for a BCFT2. Due to modified homology conditions of the RT surfaces in this framework, the holographic entanglement entropy formula involves extremal surfaces that terminate on the EOW brane Takayanagi:2011zk ; Fujita:2011fp .
@@
This AdS3/BCFT2 scenario was further developed in Deng:2020ent , where the authors considered the bulk as a defect spacetime with conformal matter on the EOW brane. In this case, the gravity region relevant to determining the QES is localized only on the EOW brane and emerges via a partial Randall-Sundrum reduction of the bulk geometry. In the lower-dimensional effective description, transparent boundary conditions were imposed at the interface of the gravitational and non-gravitational regions of the CFT2. This description, known as the defect extremal surface (DES) prescription, was proposed as a holographic counterpart of the island formula. Computations of the fine-grained entanglement entropy using the DES formula and the boundary QES prescription were shown to agree. Subsequently, the authors in Chu:2021gdb employed the DES formula to compute the entanglement entropy for subsystems in radiation baths associated with a $2d$ eternal black hole.
@@
To assess the robustness of the island framework and its double holographic generalizations, it is crucial to examine them in a less conventional gravitational setting. The solvable nature of CFTs deformed by an irrelevant operator formed from the stress-energy tensors Zamolodchikov:2004ce ; Cavaglia:2016oda ; Smirnov:2016lqw creates a suitable environment for this purpose. These irrelevant deformations, termed as $T\bar{T}$ deformations and extensively studied in the literature Shyam:2017znq ; Kraus:2018xrn ; Cottrell:2018skz ; Taylor:2018xcy ; Hartman:2018tkw ; Shyam:2018sro ; Caputa:2019pam ; Giveon:2017myj ; Asrat:2017tzd ; Donnelly:2018bef ; Lewkowycz:2019xse ; Chen:2018eqk ; Banerjee:2019ewu ; Jeong:2019ylz ; Murdia:2019fax ; Park:2018snf ; Asrat:2019end ; He:2019vzf ; Grieninger:2019zts ; Khoeini-Moghaddam:2020ymm ; Basu:2024bal ; Basu:2023aqz ; Basu:2023bov ; Basu:2024enr ; Grieninger:2023knz ; Chang:2024voo ; Basu:2024xjq ; Basu:2025exh ; Basu:2025fsf ; Hirano:2025alr ; Hirano:2025tkq ; He:2025ppz ; Pant:2024eno , introduce a finite cut-off in the AdS3 geometries McGough:2016lol with Dirichlet boundary conditions. [^1] Note that alternate holographic proposals for $T\bar{T}$ deformed theories, such as Hirano:2020nwq ; Kraus:2018xrn ; Guica:2019nzm ; Dubovsky:2018bmo ; Tolley:2019nmm ; Hirano:2025cjg , exist in the literature and are expected to coincide under certain limits.
@@
In Deng:2023pjs ; Deng:2024dct , the authors computed the entanglement entropy for bipartite pure states using both the island and the DES formula within a cut-off AdS3 spacetime truncated by an EOW brane. This setup involves two boundaries- an EOW brane characterizing the AdS/BCFT correspondence, and a finite cut-off surface that arises due to $T\bar{T}$ deformed theories. These surfaces differ fundamentally due to the distinct boundary conditions imposed and introduce subtleties in the standard AdS/BCFT correspondence, which require careful analysis. In the boundary description, this construction is equivalent to a $T\bar{T}$ deformed BCFT2 obtained via partial dimensional reduction of the braneworld gravity glued to a non-gravitating CFT with $T\bar{T}$ deformation Deng:2023pjs .
@@
In Deng:2020ent , the proposed duality between the island formula and the DES prescription has been verified using the entanglement entropy for simple braneworld models where the induced geometry on the EOW brane was Poincaré AdS2 and the BCFT was on a flat and non-gravitating background. Further verification of this duality using mixed state entanglement measures like the entanglement negativity and the reflected entropy was subsequently carried out in Basu:2022reu ; Shao:2022wrm ; Li:2021dmf , and for $T\bar{T}$ deformed scenarios in Deng:2023pjs ; Deng:2024dct ; Basu:2024xjq . However, it is crucial to verify this duality in more general settings where the induced metric on the brane and the BCFT background might not be so simple. In Wang:2021xih the authors considered a gravitational FRW cosmology on the EOW brane coupled to a flat CFT on the asymptotic boundary, and verified the equivalence between the island and the DES prescriptions using the entanglement entropy.
@@
Although significant progress has been made, a systematic test for this duality in settings where both the EOW brane and the radiation bath reside on non-trivial geometries, as well as in the presence of irrelevant deformations such as the $T\bar{T}$ deformation, is still an open problem. In this article, we address this interesting issue and investigate the proposed duality between the island and the DES prescription for braneworld models where the radiation bath and the EOW brane reside on a black hole geometry. In particular, we consider the model described in Geng:2021mic ; Geng:2022dua , where the authors consider a bulk AdS3 black string geometry truncated by a Karch-Randall (KR) brane, dual to a BCFT2 considered on an eternal AdS2 black hole background. In this model, the metric induced on the brane is also an eternal AdS2 black hole. We modify this model to include conformal matter localized on the EOW brane, and utilizing the island and the DES prescription, we compute the entanglement entropy for subsystems considered in radiation baths associated with an eternal AdS2 black hole. We show that the results obtained using the DES prescription align with those obtained from the island formula, and provide a strong verification for the proposed duality. We further extend our analysis to an AdS3 black string geometry with a finite cut-off arising from the $T\bar{T}$ deformation and truncated by the EOW brane, and compute the entanglement entropy up to first order corrections for subsystems in the radiation baths. Once again, we demonstrate an agreement between the results obtained using the bulk DES and the island prescriptions.
@@
The remainder of the article is as follows. In section˜2 we briefly discuss some necessary concepts, such as the Defect AdS3/BCFT2, the island and the DES formula for the entanglement entropy, the braneworld model of a holographic BCFT2 located on an AdS2 black hole background, and $T\bar{T}$ deformation. In section˜3, we discuss the computations of the entanglement entropy for subsystems in the radiation baths in this braneworld model, utilizing both the island and the DES prescription. Subsequently, we determine the correction to the entanglement entropy in the cut-off AdS3 black string model arising from $T\bar{T}$ deformation truncated by the EOW brane in section˜4. We generate and compare Page curves for both the undeformed and the $T\bar{T}$ deformed scenarios in section˜5, and finally, we conclude by summarizing our results and discussing their implications in section˜6.
@@ 2.1 Defect AdS3/BCFT2
In this section we briefly review the AdS3/BCFT2 correspondence first proposed in Takayanagi:2011zk ; Fujita:2011fp . A boundary conformal field theory (BCFT) is defined as a conformal field theory (CFT) with a boundary, where conformal boundary conditions are applied. The holographic dual of a BCFT Cardy:2004hm is an asymptotically AdS spacetime truncated by a co-dimension 1 end-of-the-world (EOW) brane where Neumann boundary conditions are applicable (see fig.˜1). The bulk action in gives as Takayanagi:2011zk ; Fujita:2011fp
$I_{0}=-\frac{1}{16\pi G_{N}}\int_{N}\sqrt{g}(R-2\Lambda)-\frac{1}{8\pi G_{N}}\int_{Q}\sqrt{h}(K-(d-1)T),$ (1)
@@
$K_{ab}=(K-T)h_{ab}.$ (2)
Figure 1: Holographic dual of a BCFT2 defined on a half plane $(x>0)$. The EOW brane is considered to be at a constant hyperbolic angle $\rho_{0}$. Figure modified from Deng:2020ent .
@@
$ds^{2}=d\rho^{2}+\cosh^{2}\rho\left(\frac{d\tau^{2}+dy^{2}}{y^{2}}\right),$ (3)
@@
$ds^{2}_{brane}=\Omega^{2}_{y}(d\tau^{2}+dy^{2})=\Omega^{2}_{y}\,\ ds^{2}_{flat},\qquad\Omega_{y}=\bigg|\frac{1}{y\operatorname{sech}\rho_{0}}\bigg|.$ (4)
@@
$K_{ab}=\tanh\rho_{0}\,\ h_{ab},\qquad T=\tanh\rho_{0}.$ (5)
This AdS3/BCFT2 setup was further developed in Deng:2020ent ; Chu:2021gdb where the authors incorporated conformal matter localised on a tensionless EOW brane. This induces a tension on the brane, which result in the brane to relocate to some constant angle dependent on the tension. The bulk action in this modified setup is given by
$I=I_{0}+\int_{Q}\sqrt{h}\mathcal{L}_{Q},$ (6)
where $\mathcal{L}_{Q}$ is the Lagrangian corresponding to the conformal matter on the brane. [^2] See Chen:2020uac for some examples. Once again by variation of this modified action with respect to the induced metric $h_{ab}$ and implementation of the Neumann boundary condition we arrive at Takayanagi:2011zk ; Deng:2020ent ; Chu:2021gdb
$K_{ab}=(K-T)h_{ab}+8\pi G_{N}\langle T_{ab}\rangle,\qquad\langle T_{ab}\rangle=\frac{2}{\sqrt{h}}\frac{\partial\mathcal{L}_{Q}}{\partial h_{ab}}.$ (7)
@@ 2.2 Entanglement Entropy and the DES formula
In the modified scenario of defect AdS3/BCFT2 discussed in the previous subsection, the entanglement entropy of a subsystem $A$ undergoes modifications to incorporate contributions from the conformal matter localised on the brane. In this framework the fine grained entanglement entropy of a subsystem $A$ defined in the radiation bath is computed using the island formula described in Almheiri:2019hni ; Penington:2019npb ; Almheiri:2019psf as
$S_{Is}=\text{min}\,\ \text{ext}_{X}\Big\{S_{eff}(A\cup I_{S}(A))+S_{area}(X)\Big\},\qquad X=\partial I_{S}(A)$ (8)
@@
From the holographic perspective, the Defect Extremal Surface (DES) formula was introduced as a holographic counterpart of the island formula in Deng:2020ent , where the authors enhanced the Quantum Extremal Surface (QES) formula to incorporate contributions from the bulk defect matter localized on the EOW brane. This leads to the modification of the Ryu-Takayanagi (RT) formula for the holographic entanglement entropy to the DES formula as
$S_{DES}=\text{min}\,\ \text{ext}_{\Gamma_{A},X}\Big\{S_{RT}(\Gamma_{A})+S_{defect}(D)\Big\},\qquad X=\Gamma_{A}\cap D,$ (9)
@@ 2.3 Holographic BCFT2 in a Black Hole background
In Geng:2021mic ; Geng:2022dua , the authors examined the entanglement entropy of subsystems considered on the asymptotic boundary of an AdS3 black string geometry truncated by a Karch-Randall (KR) brane. The bulk geometry is described by the metric
$ds^{2}=\frac{\cosh^{2}\rho}{u^{2}}\bigg[-\left(1-\frac{u}{u_{h}}\right)dt^{2}+\frac{du^{2}}{\left(1-\frac{u}{u_{h}}\right)}\bigg]+d\rho^{2},$ (10)
@@
$X_{2}=\frac{2\sqrt{u_{h}^{2}-uu_{h}}}{u}\cosh\frac{2\pi t}{\beta}\cosh\rho,\qquad X_{3}=\sinh\rho,$ (11)
@@
$ds^{2}=\eta_{AB}dX^{A}dX^{B},\qquad\eta_{AB}=\text{diag}(-1,-1,1,1),$ (12)
@@
$X_{A}X^{A}=-1.$ (13)
Here $\beta=4\pi u_{h}$ is the inverse Hawking temperature. Using the above embedding relations, the holographic entanglement entropy for radiation subsystems considered on both the left and right copies of the TFD state corresponding to an eternal black hole was evaluated in Geng:2022dua as
$S_{A}=\begin{cases}\frac{c}{3}(\rho_{\epsilon}+\rho_{b})&\text{Boundary channel,}\\ \frac{c}{6}\left(\log\left[\frac{u_{h}}{LR}\left(\Delta_{L}+\Delta_{R}+2\sqrt{\Delta_{L}\Delta_{R}}\cosh\frac{t}{u_{h}}\right)\right]+2\rho_{\epsilon}\right)\qquad&\text{Bulk channel,}\end{cases}$ (14)
where $\Delta_{L}=u_{h}-L$, $\Delta_{R}=u_{h}-R$, and $\rho_{\epsilon}$ is related to the UV cut-off $\epsilon$ as $\rho_{\epsilon}=\log\frac{2}{\epsilon}$. These results were also reproduced from the boundary description using BCFT correlators of twist fields in a black hole background in Geng:2022dua .